1)Parkette, Inc., acquired a 60 percent interest in Skybox Company several years ago. During 2017, Skybox sold inventory costing $270,000 to Parkette for $337,500. A total of 13 percent of this inventory was not sold to outsiders until 2018. During 2018, Skybox sold inventory costing $259,250 to Parkette for $305,000. A total of 32 percent of this inventory was not sold to outsiders until 2019. In 2018, Parkette reported cost of goods sold of $560,000 while Skybox reported $337,500. What is the consolidated cost of goods sold in 2018?

2)James Corporation owns 80 percent of Carl Corporation's common stock. During October, Carl sold merchandise to James for $240,000. At December 31, 40 percent of this merchandise remains in James's inventory. Gross profit percentages were 30 percent for James and 40 percent for Carl. The amount of intra-entity gross profit in inventory at December 31 that should be eliminated in the consolidation process is

3)Dane, Inc., owns Carlton Corporation. For the current year, Dane reports net income (without consideration of its investment in Carlton) of $216,000 and the subsidiary reports $98,500. The parent had a bond payable outstanding on January 1, with a carrying amount of $230,000. The subsidiary acquired the bond on that date for $208,500. During the current year, Dane reported interest expense of $23,760 while Carlton reported interest income of $22,860, both related to the intra-entity bond payable. What is consolidated net income?

4)Thomson Corporation owns 70 percent of the outstanding stock of Stayer, Incorporated. On January 1, 2016, Thomson acquired a building with a 10-year life for $326,000. Thomson depreciated the building on the straight-line basis assuming no salvage value. On January 1, 2018, Thomson sold this building to Stayer for $288,800. At that time, the building had a remaining life of eight years but still no expected salvage value. In preparing financial statements for 2018, how does this transfer affect the computation of consolidated net income?
Answer #1

1)

Unrealized profit of inv 2017 in 2018= (337500-270000)*0.13

=67500*0.13

=8,775

Unrealized profit of inv obtained in 2018 and sales in 2018=(305000-259250)*(1-0.32)

=45750*0.68=31,110

Cost of goods sold in skybox=337500-31,110-8775-259250

=38365

cost of goods sold in PArkette=560000

consolidated cost of goods sold in 2018=560000+38365

=598365$

2)

Unrealized Gross Profit = 240,000*40%40%

=38,400

3)

Net income of the parent 216,000
Net income of the subsidiary 98,500
Gain on retirement of bond (230,000-208,500) 21,500
Interest expense 23,760
Interest income (22,860)
Consolidated net income 336,900

4)

Depreciation charged on Building by Thomson per year = (Cost - Salvage Value)/Useful Life

= ($326,000 - $0)/10 years = $32,600 per year

Book Value of Building on the date of sale = Cost - Depreciation charged for 2016 and 2017

= $326,000 - $32,600 - $32,600 = $260,800

Profit on sale of building in the books of Thomson Corporation = Sale price - Book Value

= $288,800 - $260,800 = $28,000

This profit on sale of building is an intra-entity gain which need to be reduced from consolidated net income.

Depreciation charged by Stayer for 2018 = Sale price of Building/Estimated Useful life

= $288,800/8 years = $36,100 per year

Depreciation to be charged as building was previously owned by Thomson = $32,600 per year

Excess depreciation charged of $3,500 ($36,100 - $32,600) need to be add back to consolidated net income.

Therefore net effect of intra entity adjustments on consolidated net income will be $24,500($28,000-$3,500).

Hence, Net income is reduced by $24,500